OpenStudy (anonymous):

Solve. 8/x + 15/x^2 =-1

OpenStudy (saifoo.khan):

The LCD = x^2

OpenStudy (saifoo.khan):

\[\frac{8(x) + 15}{x^2 } = -1\]Can you solve now?

OpenStudy (anonymous):

I think so.....

OpenStudy (saifoo.khan):

Cool.

OpenStudy (anonymous):

is is x= 5, 3?

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

yay! Thank you!